Hi, I am trying to complete a practice test for my Algebra class but I need help solving the following problem:
How much water must be evaporated from 64 oz of a 10% salt solution to make a 2% salt solution?
If possible, could you also show me how you got the answer.
Thank you.
Lia M.

Answer by mananth(16946)

Firstly if you evaporate 10% solution the concentration will increase. It wont decrease.
The other way is possible
initial conc 2 % final 10%
water to be evaporated be x

when x evaporates there is 64 - x left,
which has 90% water
2% of 64 = 1.28 oz is salt,
64 - 1.28 = 62.72oz is water.

62.72 - x = 0.9 (64 - x)
62.72 - x = 57.6 - 0.90x
-0.1x = -5.12
x = 51.2 oz of water
m.ananth@hotmail.ca